# An Optimal Design Framework for Lasso Sign Recovery

### v2

@@ Abstract
Keywords: Screening experiments; constrained-positive $Var(s)$-criterion; variable selection; $UE(s^{2})$-criterion; Gauss Dantzig selector
@@ 1 Introduction
Throughout this paper, we will assume a main effects model $\boldsymbol{y}=\beta_{0}\boldsymbol{1}+\textbf{X}\boldsymbol{\beta}+\boldsymbol{e}$ where X is the $n\times p$ design matrix with elements $x_{ij}=\pm 1$, $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})^{T}$ is a sparse vector with $k<p$ nonzero elements, and $\boldsymbol{e}\sim N(\boldsymbol{0},\sigma^{2}\textbf{I})$. Without loss of generality, assume $\sigma^{2}=1$, making $\boldsymbol{\beta}$ the signed signal-to-noise ratios. Then the analysis goal is estimation of the support of $\boldsymbol{\beta}$, denoted $\mathcal{A}=\{j:|\beta_{j}|>0\}$, although for technical reasons we will be more interested in estimating the sign vector of $\boldsymbol{\beta}$, denoted $\boldsymbol{z}=\text{sign}(\beta_{j})$. We will focus our attention on optimal SSDs, although the framework developed in this paper could be applied to more general screening designs. Note that when we indicate designs are “optimal”, we are referring to the best designs found by heuristic optimality procedures, unless otherwise indicated.
@@ 1.1 Background and Open Problems
Most SSD construction methods focus on optimizing heuristic criteria that find a design whose S is close to $n\textbf{I}_{p+1}$. Booth and Cox (1962) proposed the criterion $E(s^{2})=\sum_{1<i<j\leq p}s_{ij}^{2}$ assuming balanced designs (i.e., $\textbf{X}^{T}\boldsymbol{1}=\boldsymbol{0}$), which has been extensively studied (see Lin, 1993; Wu, 1993; Nguyen, 1996; Li and Wu, 1997; Tang and Wu, 1997; Ryan and Bulutoglu, 2007). Jones and Majumdar (2014) proposed the unconditional $E(s^{2})$-criterion $UE(s^{2})=\sum_{0\leq i<j\leq p}s_{ij}^{2}$ that includes the $s_{0j}^{2}$ elements. Other criteria have been proposed that measure a design’s quality in terms of subsets of columns (Deng et al., 1996, 1999; Jones et al., 2009), which is connected to a stepwise or all-subsets analysis (Lin, 1995; Abraham et al., 1999; Westfall et al., 1998; Liu et al., 2007). Chen and Lin (1998) and Sarkar et al. (2009) investigated support recovery properties of SSDs under such sequential analyses.
Li and Lin (2002) and Phoa et al. (2009) advocated for SSDs to be analyzed under penalized estimation that induces sparse estimates. The previously mentioned SSD criteria, however, are not directly tied to statistical properties of such an analysis. There have been some attempts in the literature to address this disconnect. Marley and Woods (2010), Draguljić et al. (2014), and Weese et al. (2015) have compared optimal SSDs under different criteria via simulation with the lasso and/or Gauss-Dantzig selector (GDS), finding little difference between the SSDs in terms of support recovery. Recently, Weese et al. (2017) and Weese et al. (2021) proposed $Var(s+)$-optimal designs that minimize
$Var(s)=UE(s^{2})-UE(s)^{2}\text{ such that }\frac{UE^{*}(s^{2})}{UE(s^{2})}>c\text{ and }UE(s)>0\ ,$
where $UE^{*}(s^{2})$ is the optimal $UE(s^{2})$ value for the given $n$ and $p$, and $UE(s)=\sum_{0\leq i<j\leq p}s_{ij}$. The ideal structure of S under $Var(s+)$ is notably non-orthogonal, having average off-diagonal elements equal to a small, positive constant, making $Var(s)=0$ while satisfying the constraints on $UE(s^{2})$ and $UE(s)$. Simulation studies showed $Var(s+)$-optimal designs (with $c=0.8$) performed similarly to other optimal SSDs under the GDS, but when the effect directions were known in advance (i.e. the nonzero signs of $\boldsymbol{\beta}$ were all positive), the $Var(s+)$-optimal design had better support recovery properties. The simulation studies mentioned above only perform relative comparisons of SSDs generated by heuristic criteria and so do not answer the following open problems:
@@
    Why does the $Var(s+)$-criterion identify SSDs with better support recovery probability when the active effects have positive signs?
@@
Ideally, EC would be 0, corresponding to an orthogonal design. Their second criterion assumes the sign of the active effects is always positive, denoted by $\boldsymbol{1}_{k}$. First, for a given $\mathcal{A}$, the following two weak sign consistency conditions are checked
$|\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}_{k}|$ $<\boldsymbol{1}_{p-k}$ (1)
@@
where $\textbf{C}_{\mathcal{I}\mathcal{A}}$ is the matrix of cross correlations of columns between the inactive set, =$\mathcal{I}=\{j:\beta_{j}=0\}$, and $\mathcal{A}$. Similar to EC, this condition is checked across all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\gamma_{k}$ denotes the proportion of such $\mathcal{A}$ that achieve both conditions. Finally, their weak sign consistency criterion is defined to be
@@
We want to find a design that maximizes WSC. It would be preferred that the left hand sides of (1) and (2) equal 0, again corresponding to orthogonal designs, but nonorthogonal designs may still have a large WSC. The authors employ a hybrid elitist Pareto optimality-based coordinate exchange algorithm (Cao et al., 2017) to identify designs on the Pareto front of the EC and WSC. We refer to their optimal designs as Dantzig-based Consistency Designs (DCDs).
@@
The DCD criteria require evaluations across all $\mathcal{A}$ of sizes $k=2,\dots,\lceil n/2\rceil$. To ease the computational burden, Singh and Stufken (2022) randomly sample subsets for each $k$. Furthermore, their examples initialized their search algorithms with only two designs each, a $Var(s+)$-optimal design and either a random, $E(s^{2})$-optimal, or $UE(s^{2})$-optimal design. The resulting DCDs initialized at a $Var(s+)$ design were shown via simulation to have better support recovery properties than the $Var(s+)$-optimal design assuming positive signs of the active effects. However, their improvement over the $Var(s+)$-optimal design may be attributed to their initialization at that design. That is, they show that the initialized $Var(s+)$-optimal design is not globally optimal for known signs, but there is no guarantee a DCD is either. Indeed, an orthogonal design is the ideal design under their criteria. It may be that the globally optimal DCD performs worse than the $Var(s+)$ design. Still, their results provide strong support for constructing SSDs under criteria directly tied to statistical properties of penalized estimators.
@@
Others have considered the role of X in the statistical properties of the lasso, though not in the context of SSDs. Zhao and Yu (2006) identified the strong irrepresentable condition (SIC) on X for establishing support recovery consistency of the lasso (Zhang and Huang, 2008; Jia and Rohe, 2015). Wainwright (2009) and Hastie et al. (2019) have studied support recovery for random designs where the rows of X are independently generated from a zero-mean multivariate Normal distribution. Random designs, however, generally perform poorly for small run designs like SSDs. Based on the lasso’s SIC, Deng et al. (2013) constructed SSDs from nearly orthogonal Latin hypercube designs to minimize the off-diagonals of S assuming factors have settings between $[0,1]$. This is essentially a construction technique for minimizing $UE(s^{2})$ with more general factor settings, even though SSDs commonly assume fixed settings of ±$\pm 1$. Xing (2015) proposed a lasso SIC criterion to construct optimal two-level SSDs, similar to the $\gamma_{k}$ from (3). Huang et al. (2020) proposed a lasso optimal design theory that applies variance-based criteria to the approximate covariance matrix of the debiased lasso (Javanmard and Montanari, 2014), which is capable of performing inference via confidence intervals. Under the framework of approximate designs (i.e., $n\to\infty$) they note their criteria are not convex and give an equivalence theorem for establishing whether a design is locally optimal. They then propose an algorithmic construction for generating many local optimal approximate designs and implement a rounding procedure on the approximate design’s replication weights to produce an exact design. Overall, their approach requires many stages of approximation, leading to a discrepancy between the approximate and exact design’s covariance matrix.
@@ 1.2 Contributions and Overview
This paper addresses the two SSD open problems by developing optimality criteria that directly maximize the probability of sign recovery for the lasso estimator. The Dantzig selector and its statistical properties are closely related to the lasso estimator (Meinshausen et al., 2007; Lounici, 2008; Bickel et al., 2009; James et al., 2009; Asif and Romberg, 2010). The statistical properties of the lasso are tractable, making its optimal design framework more straightforward.
The first major contribution of this paper is a local optimal design approach to maximize the sign recovery probability assuming $\boldsymbol{\beta}$ is known and the lasso tuning parameter, $\lambda>0$, is fixed. We then develop criteria that relax the model assumptions and summarize the sign recovery probability across a range of $\lambda$. As optimization of the proposed criteria is challenging, our second major contribution is a criterion based on the $p\times p$ correlation matrix of X. We show the ideal correlation matrices under different heuristic criteria are optimal under our new criteria, thus providing solutions to both open problems. In particular, we prove the surprising result that an orthogonal correlation matrix is suboptimal when $\boldsymbol{z}$ is known. Our final contribution is a computationally efficient construction algorithm that leverages the speed of exchange algorithms for optimizing heuristic criteria.
The paper is organized as follows. Section 2 develops exact design optimality criteria targeting the lasso’s sign recovery probability under different assumptions about the model. Section 3 presents a pseudo-approximate optimal design framework that targets the optimal correlation matrix of X and justifies the ideal designs sought after by heuristic criteria under different assumptions about $\boldsymbol{z}$. Section 4 describes a computationally efficient algorithm for constructing exact optimal designs under our proposed criteria. We then demonstrate the optimal SSDs under our new framework in Section 5. We conclude the paper with a discussion in Section 6, describing important implications of our results and future work.
@@ 2 Exact Local Optimality Criteria
where $\textbf{C}=n^{-1}\textbf{F}^{T}\textbf{F}$ is the correlation matrix of the columns of X. Denote the estimated support and sign of the resulting lasso estimate, $\hat{\boldsymbol{\beta}}$, by $\hat{\mathcal{A}}=\{j:|\hat{\beta}_{j}|>0\}$ and $\hat{\boldsymbol{z}}=\text{sign}(\hat{\beta}_{j})$, respectively.
@@
where $\textbf{C}_{\hat{\mathcal{A}}}^{-1}=n(\textbf{F}_{\hat{\mathcal{A}}}^{T}\textbf{F}_{\hat{\mathcal{A}}})^{-1}$, $\hat{\textbf{Z}}_{\hat{\mathcal{A}}}=\text{Diag}(\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}})$, and ^=$\hat{\mathcal{I}}=\{j:\hat{\beta}_{j}=0\}$. For given X, $\boldsymbol{y}$, and $\boldsymbol{\beta}$, we can check whether the resulting $\hat{\boldsymbol{\beta}}^{*}$ has the true support and sign vector by setting $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$ and checking the KKT conditions. If they hold, $\hat{\boldsymbol{\beta}}^{*}$ recovers the sign, which is more stringent than recovering the support. The proposed criteria in this section rank designs according to the probability of sign recovery: $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})$, which depends on the unknown parameter values and $\lambda$. Therefore, to calculate the probability, one must specify a $\lambda>0$ and a $\boldsymbol{\beta}$, so our proposed framework falls in the class of local optimal designs commonly employed with nonlinear models and maximum likelihood estimation (Silvey, 1980; Khuri et al., 2006; Yang and Stufken, 2009). This should not be particularly surprising because the probability is similar to a power calculation, which requires specification of effect sizes. We will recall this connection in Section 2.2 where we introduce more practical criteria.
Clearly $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})$ is a joint probability of two events. The first event follows from equations (4) and (5) and checks whether $\hat{\boldsymbol{\beta}}_{\mathcal{A}}$ has the sign $\boldsymbol{z}_{\mathcal{A}}$:
$S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}}$ =$=\left\{\boldsymbol{u}<\sqrt{n}\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}\right\}\ ,$ (7)
where $\boldsymbol{u}\sim N(\lambda\sqrt{n}\,\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{Z}_{\mathcal{A}})$. Note we define the event with respect to $\boldsymbol{\beta}_{\mathcal{A}}$ rather than the design-dependent $\boldsymbol{\beta}_{\mathcal{A}}^{*}$. The second event follows from equations (4) and (6) and coincides with all $j\in\mathcal{I}$ having $\hat{\beta}^{*}_{j}=0$:
$I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{z}=\left\{|\boldsymbol{v}|\leq\lambda\sqrt{n}\boldsymbol{1}\right\}\ ,$ (8)
where $\boldsymbol{v}~{}N(\lambda\sqrt{n}\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{C}_{\mathcal{I}}-\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{C}_{\mathcal{A}\mathcal{I}})$. The event depends on $\boldsymbol{\beta}$ only though $\boldsymbol{z}$.

### v3

@@ Abstract
Keywords: Constrained-positive $\text{Var}(s)$-criterion; Gauss Dantzig selector; Supersaturated designs; $\text{UE}(s^{2})$-criterion; Variable selection
@@ 1 Introduction
Throughout this paper, we will assume a main effects model $\boldsymbol{y}=\beta_{0}\boldsymbol{1}+\textbf{X}\boldsymbol{\beta}+\boldsymbol{e}$ where X is the $n\times p$ design matrix with elements $x_{ij}=\pm 1$, $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})^{T}$ is a sparse vector with $k<p$ nonzero elements, and $\boldsymbol{e}\sim N(\boldsymbol{0},\sigma^{2}\textbf{I})$. Without loss of generality, assume $\sigma^{2}=1$, making $\boldsymbol{\beta}$ the signed signal-to-noise ratios. Then the analysis goal is estimation of the support of $\boldsymbol{\beta}$, denoted $\mathcal{A}=\{j:|\beta_{j}|>0\}$, although for technical reasons we will be more interested in estimating the sign vector of $\boldsymbol{\beta}$, denoted $\boldsymbol{z}=\text{sign}(\boldsymbol{\beta})$. We will focus our attention on optimal SSDs, although the framework developed in this paper could be applied to more general screening designs. Note that when we indicate designs are “optimal”, we are referring to the best designs found by heuristic optimality procedures, unless otherwise indicated.
@@ 1.1 Background and Open Problems
Most SSD construction methods focus on optimizing heuristic criteria that find a design whose S is close to $n\textbf{I}_{p+1}$. Booth and Cox (1962) proposed the criterion $\text{E}(s^{2})=2/(p(p-1))\sum_{1<i<j\leq p}s_{ij}^{2}$ assuming balanced designs (i.e., $\textbf{X}^{T}\boldsymbol{1}=\boldsymbol{0}$), which has been extensively studied (see Lin, 1993; Wu, 1993; Nguyen, 1996; Li and Wu, 1997; Tang and Wu, 1997; Ryan and Bulutoglu, 2007). Jones and Majumdar (2014) proposed the unconditional $\text{E}(s^{2})$-criterion, $\text{UE}(s^{2})=2/(p(p+1))\sum_{0\leq i<j\leq p}s_{ij}^{2}$, that includes the $s_{0j}^{2}$ elements. Other criteria have been proposed that measure a design’s quality in terms of subsets of columns (Deng et al., 1996, 1999; Jones et al., 2009), which is connected to a stepwise or all-subsets analysis (Lin, 1995; Abraham et al., 1999; Westfall et al., 1998; Liu et al., 2007). Chen and Lin (1998) and Sarkar et al. (2009) investigated support recovery properties of SSDs under such sequential analyses.
Li and Lin (2002) and Phoa et al. (2009) advocated for SSDs to be analyzed under penalized estimation that induces sparse estimates. The previously mentioned SSD criteria, however, are not directly tied to statistical properties of such an analysis. There have been some attempts in the literature to address this disconnect. Marley and Woods (2010), Draguljić et al. (2014), and Weese et al. (2015) have compared optimal SSDs under different criteria via simulation with the lasso and/or Gauss-Dantzig selector (GDS), finding little difference between the SSDs in terms of support recovery. Additionally, the Dantzig selector and its statistical properties are closely related to the lasso estimator (Meinshausen et al., 2007; Lounici, 2008; Bickel et al., 2009; James et al., 2009; Asif and Romberg, 2010), but the mathematics of the lasso are more tractable, making its optimal design framework more straightforward.
Recently, Weese et al. (2017) and Weese et al. (2021) proposed $\text{Var}(s+)$-optimal designs that minimize
$\text{Var}(s)=\text{UE}(s^{2})-\text{UE}(s)^{2}\text{ such that }\frac{\text{UE}^{*}(s^{2})}{\text{UE}(s^{2})}>c\text{ and }\text{UE}(s)>0\ ,$
where $\text{UE}^{*}(s^{2})$ is the optimal $\text{UE}(s^{2})$ value for the given $n$ and $p$, and $\text{UE}(s)=2/(p(p+1))\sum_{0\leq i<j\leq p}s_{ij}$. The ideal structure of S under $\text{Var}(s+)$ is notably non-orthogonal, having average off-diagonal elements equal to a small, positive constant, making $\text{Var}(s)=0$ while satisfying the constraints on $\text{UE}(s^{2})$ and $\text{UE}(s)$. Simulation studies showed $\text{Var}(s+)$-optimal designs (with $c=0.8$) performed similarly to other optimal SSDs under the GDS, but when the effect directions were known in advance (i.e. the nonzero signs of $\boldsymbol{\beta}$ were all positive), the $\text{Var}(s+)$-optimal design had better support recovery properties. The simulation studies mentioned above only perform relative comparisons of SSDs generated by heuristic criteria and so do not answer the following open problems:
@@
    Why does the $\text{Var}(s+)$-criterion identify SSDs with better support recovery probability when the active effects have positive signs?
@@
Ideally, EC would be 0, corresponding to an orthogonal design. Their second criterion assumes the sign of the active effects is always positive, denoted by $\boldsymbol{1}_{k}$. For a given $\mathcal{A}$, the following two weak sign consistency conditions are checked
$|\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}_{k}|$ $<\boldsymbol{1}_{p-k}\ ,$ (1)
@@
where $\textbf{C}_{\mathcal{I}\mathcal{A}}$ is the matrix of cross correlations of columns between the inactive set, =$\mathcal{I}=\{j:\beta_{j}=0\}$, and $\mathcal{A}$. Similar to EC, this condition is checked across all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\gamma_{k}$ denotes the proportion of such $\mathcal{A}$ that achieve both (1) and (2). Their weak sign consistency criterion is defined to be
@@
A design that maximizes WSC is then highly likely to satisfy the two weak sign consistency conditions across many $\mathcal{A}$ where $|\mathcal{A}|\leq\lceil n/2\rceil$. It would be preferred that the left hand sides of (1) and (2) equal 0, again corresponding to orthogonal designs, but nonorthogonal designs may still have a large WSC. The authors employ a hybrid elitist Pareto optimality-based coordinate exchange algorithm (Cao et al., 2017) to identify designs on the Pareto front of the EC and WSC. We refer to their optimal designs as Dantzig-based Consistency Designs (DCDs).
@@
The DCD criteria require evaluations across all $\mathcal{A}$ of sizes $k=2,\dots,\lceil n/2\rceil$. To ease the computational burden, Singh and Stufken (2023) randomly sample subsets for each $k$. Furthermore, their examples initialized their search algorithms with only two designs each, a $\text{Var}(s+)$-optimal design and either a random, $\text{E}(s^{2})$-optimal, or $\text{UE}(s^{2})$-optimal design. The resulting DCDs initialized at a $\text{Var}(s+)$ design were shown via simulation to have better support recovery properties than the $\text{Var}(s+)$-optimal design assuming positive signs of the active effects. However, their improvement over the $\text{Var}(s+)$-optimal design may be attributed to their initialization at that design. That is, they show that the initialized $\text{Var}(s+)$-optimal design is not globally optimal for known signs, but there is no guarantee a DCD is either. Indeed, an orthogonal design is the ideal design under their criteria. It may be that the globally optimal DCD performs worse than the $\text{Var}(s+)$ design. Still, their results provide strong support for constructing SSDs under criteria directly tied to statistical properties of penalized estimators.
@@
Others have considered the role of X in the statistical properties of the lasso, though not in the context of SSDs. Zhao and Yu (2006) identified the strong irrepresentable condition (SIC) on X for establishing support recovery consistency of the lasso (Zhang and Huang, 2008; Jia and Rohe, 2015). Wainwright (2009) and Hastie et al. (2019) have studied support recovery for random designs where the rows of X are independently generated from a zero-mean multivariate Normal distribution. Random designs, however, generally perform poorly for small run designs like SSDs. Based on the lasso’s SIC, Deng et al. (2013) constructed SSDs from nearly orthogonal Latin hypercube designs to minimize the off-diagonals of S assuming factors have settings between $[0,1]$. This is essentially a construction technique for minimizing $\text{UE}(s^{2})$ with more general factor settings, even though SSDs commonly assume fixed settings of ±$\pm 1$. Xing (2015) proposed a lasso SIC criterion to construct optimal two-level SSDs, similar to the $\gamma_{k}$ from (3). Huang et al. (2020) proposed a lasso optimal design theory that applies variance-based criteria to the approximate covariance matrix of the debiased lasso (Javanmard and Montanari, 2014), which is capable of performing inference via confidence intervals. Under the framework of approximate designs (i.e., $n\to\infty$) they note their criteria are not convex and give an equivalence theorem for establishing whether a design is locally optimal. They then propose an algorithmic construction for generating many local optimal approximate designs and implement a rounding procedure on the approximate design’s replication weights to produce an exact design. Overall, their approach requires many stages of approximation, leading to a discrepancy between the approximate and exact design’s covariance matrix.
@@ 1.2 Contributions and Overview
This paper addresses the two SSD open problems by developing optimality criteria that directly maximize the probability of sign recovery for the lasso estimator.
The first major contribution of this paper is a local optimal design approach to maximize the sign recovery probability assuming $\boldsymbol{\beta}$ is known and the lasso tuning parameter, $\lambda>0$, is fixed. We then develop criteria that relax the model assumptions and summarize the sign recovery probability across a range of $\lambda$. As optimization of the proposed criteria is challenging, the second major contribution is a criterion based on the $p\times p$ correlation matrix of X. We provide evidence that the ideal correlation matrices under different heuristic criteria are optimal under the new criteria, thus making substantial progress on both open problems. In particular, we prove the surprising result that an orthogonal correlation matrix is suboptimal when $\boldsymbol{z}$ is known. The final contribution is a computationally efficient construction algorithm that leverages the speed of exchange algorithms for optimizing heuristic criteria.
The paper is organized as follows. Section 2 develops exact design optimality criteria targeting the lasso’s sign recovery probability under different assumptions about the model. Section 3 presents a pseudo-approximate optimal design framework that targets the optimal correlation matrix of X and justifies the ideal designs sought after by heuristic criteria under different assumptions about $\boldsymbol{z}$. Section 4 describes a computationally efficient algorithm for constructing exact optimal designs under the proposed criteria. We then demonstrate the optimal SSDs under the new framework in Section 5. We conclude the paper with a discussion in Section 6, describing important implications of the results and future work.
@@ 2 Exact Local Optimality Criteria
where $\textbf{C}=n^{-1}\textbf{F}^{T}\textbf{F}$ is the correlation matrix of the columns of X. Denote the estimated support and sign of the resulting lasso estimate, $\hat{\boldsymbol{\beta}}^{*}$, by $\hat{\mathcal{A}}=\{j:|\hat{\beta}_{j}^{*}|>0\}$ and $\hat{\boldsymbol{\boldsymbol{z}}}=\text{sign}(\hat{\boldsymbol{\beta}}^{*})$, respectively.
@@
where $\textbf{C}_{\hat{\mathcal{A}}}^{-1}=n(\textbf{F}_{\hat{\mathcal{A}}}^{T}\textbf{F}_{\hat{\mathcal{A}}})^{-1}$, $\hat{\textbf{Z}}_{\hat{\mathcal{A}}}=\text{Diag}(\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}})$, and ^=$\hat{\mathcal{I}}=\{j:\hat{\beta}_{j}^{*}=0\}$. For given X, $\boldsymbol{y}$, and $\boldsymbol{\beta}$, we can check whether the resulting $\hat{\boldsymbol{\beta}}^{*}$ has the true support and sign vector by setting $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$ and checking the KKT conditions. If they hold, $\hat{\boldsymbol{\beta}}^{*}$ recovers the sign, which is more stringent than recovering the support.
The proposed criteria in this section rank designs according to the probability of sign recovery, denoted $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$. To calculate the probability, one must specify a $\lambda>0$ and a $\boldsymbol{\beta}$, so the proposed framework resembles optimality frameworks employed for nonlinear models and maximum likelihood estimation (Silvey, 1980; Khuri et al., 2006; Yang and Stufken, 2009). Note that specifying $\boldsymbol{\beta}$ also leads to specifying $\mathcal{A}$, $\mathcal{I}$, and $\boldsymbol{z}$. The criterion $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta},\lambda)$ is a joint probability of two events corresponding to (5) and (6) assuming $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$. The first event corresponds to (5) and checks whether $\text{sign}(\hat{\boldsymbol{\beta}}^{*}_{\mathcal{A}})=\boldsymbol{z}_{\mathcal{A}}$ where $\hat{\boldsymbol{\beta}}^{*}_{\mathcal{A}}$ is from (4) assuming $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$:
$S_{\lambda}$$\coloneqq\left\{\boldsymbol{u}<\sqrt{n}\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}\right\}\ ,$ (7)
where $\boldsymbol{u}\sim N(\lambda\sqrt{n}\,\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{Z}_{\mathcal{A}})$. The second event corresponds to (6), which checks whether all $j\in\mathcal{I}$ have $\hat{\beta}^{*}_{j}=0$:
$I_{\lambda}\coloneqq\left\{|\boldsymbol{v}|\leq\lambda\sqrt{n}\boldsymbol{1}\right\}\ ,$ (8)
where $\boldsymbol{v}\sim N(\lambda\sqrt{n}\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{C}_{\mathcal{I}}-\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{C}_{\mathcal{A}\mathcal{I}})$. The probability of event (8) depends on $\boldsymbol{\beta}$ only through $\boldsymbol{z}$. The probabilities of both events depend on X only through its C and V, a fact we exploit in Section 3. For brevity, we will sometimes refer to these events as simply $S_{\lambda}$ and $I_{\lambda}$.